Q;- Sugest two transitions in the atomic spectrum of hydrogen for which the wave no ratio is 108:7 ?


ans:- first lyman series and first pachen series


 


what is its concept pls tell

Grade:12th Pass

1 Answers

rajula vineet reddy
12 years ago

the ritz-combination equation also called as rydbergs equation is the concept

the equation is used in calculating the wave length,frequency  and wave number.

\frac{1}{\lambda_{vac}} = R_\infty \left(\frac{1}{n_1^2}-\frac{1}{n_2^2}\right)

Where

\lambda_{vac} \! is the wavelength of electromagnetic radiation emitted in vacuum,
R_\infty \! is the Rydberg constant, approximately 1.097 * 107 m-1,
n_1\! and n_2\! are integers such that n_1 < n_2\!.

By setting n1 to 1 and letting n2 run from 2 to infinity, the spectral lines known as the Lyman series converging to 91 nm are obtained, in the same manner:

n1n2NameConverge toward
1 2 → ∞ Lyman series ??91.13 nm (UV)
2 3 → ∞ Balmer series ?364.51 nm (Visible)
3 4 → ∞ Paschen series ?820.14 nm (IR)
4 5 → ∞ Brackett series 1458.03 nm (IR)
5 6 → ∞ Pfund series 2278.17 nm (IR)
6 7 → ∞ Humphreys series

3280.56 nm (IR)    

   this is the concept                                                       ---------vineet